OpenStudy (anonymous):

In solving the equation (x + 3)(x – 3) = -8, Eric stated that the solution would be x + 3 = -8 => x = -11 or (x – 3) = -8 => x = -5 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

Remember: \[(a+b)(a-b) = a^2 - b^2\] Using this: \[x^2 - 9 = -8\] Or, \[x^2 = 1\] Here you can find x and that will be your answer...

OpenStudy (anonymous):

thnaks

OpenStudy (anonymous):

Welcome dear...